OpenStudy (anonymous):

the hands of a clock have lengths 3cm and 5cm. Find the distance between the tips of the hands at 4 o´clock. Worked solutions please

OpenStudy (mertsj):



OpenStudy (mertsj):

The angle made by the hands is 120 degrees. How about the law of cosines to find the missing side of the triangle?

OpenStudy (anonymous):

how do i measure that angle?

OpenStudy (mertsj):

\[c^2=a^2+b^2-2bc \cos C\]

OpenStudy (mertsj):

How many degrees in a circle?

OpenStudy (anonymous):

360

OpenStudy (mertsj):

Since there are 12 hours on the clock face, divide 360 by 12.

OpenStudy (mertsj):

Since there are 12 hours on the clock face, divide 360 by 12.

OpenStudy (anonymous):

thaanks! :)

OpenStudy (mertsj):

Do you understand?

OpenStudy (mertsj):

When the hand moves from the 12 to the 1, it goes 30 degrees. So when it goes from the 12 to the 4 it goes 4 times 30 or 120 degrees.

OpenStudy (anonymous):

yes! thank you!

OpenStudy (mertsj):

yw